OpenStudy (anonymous):

The equation of a circle is (x + 6)2 + (y - 4)2 = 16. The point (-6, 8) is on the circle. What is the equation of the line that is tangent to the circle at (-6, 8)?

OpenStudy (ivancsc1996):

Know calculus?

OpenStudy (anonymous):

noo

OpenStudy (ankit042):

I think no need for calculus.....can you find the center of circle?

OpenStudy (anonymous):

no

OpenStudy (ivancsc1996):

Ok, you are right, I see how you are going to solve.

OpenStudy (ankit042):

(x-a)^2+(y-b)^2 =r^2 is the standard equation of circle with center at (a,b) and radius r Now can you figure out value of (a,b)

OpenStudy (ankit042):

@heygurl123 dere?

OpenStudy (anonymous):

huh?

OpenStudy (ankit042):

can you find the center of circle now?

OpenStudy (anonymous):

still no

OpenStudy (ankit042):

(x-a)^2+(y-b)^2 =r^2 is the standard equation (x + 6)^2 + (y - 4)^2 = 16 in the question can be written as (x -(- 6))^2 + (y -( 4))^2 = 4^2

OpenStudy (ankit042):

I think now you can find value of (a,b)

OpenStudy (dumbcow):

@ankit042 made it pretty clear for every circle of form \[(x-a)^{2} +(y-b)^{2} = r^{2}\] center is point (a,b) radius is r you are given \[(x+6)^{2} + (y-4)^{2} = 16\] x+6 = x-a --> a = -6 y-4 = y-b --> b=4 r^2 = 16 r = 4